OpenStudy (anonymous):

The probability of event X is 0.4. If events X and Y are complements, what is the probability of event Y?

OpenStudy (anonymous):

0.6

OpenStudy (callisto):

Since events X and Y are complements, P(X) +P(Y) =1 0.4 + P(Y) =1 P(Y) = 1-0.4 = ??

OpenStudy (anonymous):

0.6